OpenStudy (somegirl1999):

Find the length of the side of a square with an area of 121. (1 point) 484 30.25 11 6.5

OpenStudy (anonymous):

well a square - you know that both l and w will be the same so take the square root of 121

OpenStudy (nubeer):

length of one side =x area of square would be = x*x 121 =x^2 just take square root on both sides

OpenStudy (somegirl1999):

ok thanks guys

OpenStudy (anonymous):

11 = length of side

OpenStudy (somegirl1999):

thank you :P